OpenStudy (natsent06):

Select the TRUE statement. A) 15 + 13 = 15 - 13 B) 13 + 15 = 13 - 15 C) Both A and B D) Neither A nor B

OpenStudy (igreen):

Can't you tell? A: 15 + 13 = 15 - 13 28 = 2 Incorrect. B: 13 + 15 = 13 - 15 28 = -2 Incorrect. Your correct option will be D, because neither Option A or Option B are correct.

OpenStudy (natsent06):

I all ready knew that but i wanted to make sure